OpenStudy (drummergirl7):

OMG PLZ help My math is HARD and I'm really bad at fractions PLZ HELP FAST!! I'll medal for whoever is right! R/12 = 1 9/10 A. 22 4/5 B. 12 9/10 C. 10 4/5 D. 19/120 pizza help FAST

OpenStudy (drummergirl7):

Omg plz not pizza lol

OpenStudy (benlindquist):

What is ur answer

OpenStudy (drummergirl7):

Honestly I have no idea! :( I'm an idiot lol

OpenStudy (benlindquist):

take a guess

OpenStudy (drummergirl7):

Maybe B??

OpenStudy (anonymous):

I can help

OpenStudy (drummergirl7):

Yay! :)

OpenStudy (anonymous):

its c

OpenStudy (kxttxn):

1 9/10 can also be 19/10, which should make it easier to solve. Do you know how to cross multiply?

OpenStudy (drummergirl7):

I need all the help I can get lol

OpenStudy (drummergirl7):

Yay thank u so much! Can I guys help me with some more plz

OpenStudy (benlindquist):

yea

OpenStudy (kxttxn):

It's better to help her solve than to give out answers.. that way she can answer similar questions in the future.

OpenStudy (drummergirl7):

Solve. 36 = 5/9k A. 4/5 B. 20 C. 64 4/5 D. 180

OpenStudy (drummergirl7):

I'm in k12 bye

OpenStudy (drummergirl7):

And I need to do 6 more maths till 6:00 or my mom is gonna ground me for a month so I needs help lol I am down on my % but my grades are good

OpenStudy (drummergirl7):

Typo not bye I meant BTW

OpenStudy (anonymous):

do u need help with that last question?

OpenStudy (drummergirl7):

Yes! :)

OpenStudy (anonymous):

ok hold on a sec...

OpenStudy (drummergirl7):

K I just need answers fast lol 😂 I don't wanna be grounded for Christmas

OpenStudy (anonymous):

i believe its a

OpenStudy (drummergirl7):

Ok thank u

OpenStudy (anonymous):

ur welcome :)

OpenStudy (drummergirl7):

Could u help me with one more plzzzzzzzz

OpenStudy (benlindquist):

yea

OpenStudy (anonymous):

yes

OpenStudy (drummergirl7):

Which equation can be used to solve this problem? Juanita hiked 5 1/4 miles on Saturday. That was 3/4 as far as Laura Hiked. How far did Laura Hike?? A. m = 3/4 × 5 1/4 B. 3/4 m = 5 1/4 C. 3/4 ÷ m = 5 1/4 D. m- 3/4 =5 1/4

OpenStudy (drummergirl7):

Omg sorry guys I'm back my dad walked in and I had to take my dog out lol sorry

OpenStudy (drummergirl7):

@benlindquist

OpenStudy (drummergirl7):

@celestials101

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

also ITS ME! and yes

OpenStudy (drummergirl7):

Wait ur Ben?!

OpenStudy (anonymous):

-_- im golden freddy

OpenStudy (drummergirl7):

Lol 😂 I mean are u like Ben or cele? 😂😂

OpenStudy (drummergirl7):

Btw at target they have fnaf plushie key chains 😂

OpenStudy (anonymous):

i dont even know who your talking about

OpenStudy (drummergirl7):

Lol nvm 😂😂😂😂😂

OpenStudy (anonymous):

golden freddy make this sound lhgfzlnfxmhgbnmkfkm.nvdf.kfdx

OpenStudy (drummergirl7):

Omg hahahhaha IKR

OpenStudy (drummergirl7):

He be like grahhhhggggahhh

OpenStudy (drummergirl7):

And then baloonboy be like hello.....

OpenStudy (drummergirl7):

Lol

OpenStudy (anonymous):

and ben as in bendrowned

OpenStudy (anonymous):

brb

OpenStudy (drummergirl7):

Lol ok

OpenStudy (drummergirl7):

AW MAN COMMON I FAILED MY TEST DANG IT

OpenStudy (drummergirl7):

Megax I were the only one right 😂

OpenStudy (drummergirl7):

Typo U were the only one right

OpenStudy (anonymous):

lol inser mlg noises here

OpenStudy (anonymous):

do you know what 2.5 is in a fraction

OpenStudy (drummergirl7):

Here's the answers for anyone that needs help on this test 1. Is a 34 I didn't put that question btw cuz my cousin answered it 2. C 3. C 4. B 5. D

OpenStudy (drummergirl7):

No lol I'm REALLY BAD AT THEM it's like I'm 5 lol

OpenStudy (drummergirl7):

I got 6% -_-

OpenStudy (drummergirl7):

Typo 60%

OpenStudy (anonymous):

wow hay to be fair i may or may not have enlisted the internet to do that for me

OpenStudy (drummergirl7):

Lol I do that sometimes lol 😂

OpenStudy (drummergirl7):

Hey I gtg thanks for the help Megax! Lol bye